Circular No. 8802

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 XG_{16} (SPACEWATCH) C. W. Hergenrother, University of Arizona, informs us that R-band

images taken with the Catalina 1.54-m telescope show the Spacewatch discovery 2006 XG_{16} (cf. MPS 192606, 195581; MPO 113007; discovery observation below) to be a comet. It was suspected to be cometary on images taken on 2007 Jan. 27.19 UT in poor seeing. Co-added 600-s R-band exposures in excellent seeing on Jan. 28.19 show the object at R = 18.2 with a coma 9′′ in diameter and a tail 20′′ long in p.a. 135^{o}.

2006 UT  α_{2000}  δ_{2000}
Dec. 10.41280  7^{h}31^{m}13.^{s}50  +26^{o}37′28′′.7

The following orbital elements (together with the Jan. 27–28 astrometry, full residuals, and an ephemeris) are given on MPEC 2007-B79.

Epoch = 2007 Mar. 1.0 TT

T = 2007 Feb. 9.9575 TT  ω = 41.^{o}3095
e = 0.420981  Ω = 78.4544 }2000.0
q = 2.102218 AU  i = 9.0781

a = 3.630654 AU n^{o} = 0.1424711 P = 6.92 years

SATELLITES OF NEPTUNE Further to IAUC 5347, the IAU Working Group for Planetary System

Nomenclature has approved the following new designations and names of satellites of Neptune:

Neptune IX Halimede = S/2002 N 1
Neptune X  Psamathe = S/2003 N 1
Neptune XI Sao  = S/2002 N 2
Neptune XII Laomedeia = S/2002 N 3
Neptune XIII Neso  = S/2002 N 4

COMET C/2006 P1 (McNAUGHT) Visual total-magnitude and tail-length estimates (cf. IAUC 8797): Jan.

22.51 UT, −0.3, 24^{o} (A. Pearce, City Beach, W. Australia, naked eye); 23.53, +1.5, 30^{o} (D. J. Frew, Perth Observatory, naked eye); 24.53, −0.3, 26^{o} (Pearce); 25.52, +0.4, 20^{o} (Pearce); 29.65, +2.3:, 15^{o} (J. Drummond, Gisborne, New Zealand, naked eye); 31.52, 2.5, > 7^{o} (Frew); Feb. 3.95, 3.0,

– (W. Souza, São Paulo, Brazil, 8×30 binoculars).

2007 February 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green